Muffins are delivered daily to Starbucks at Science Complex. Demand is uniformly distributed between 35 and 80 muffins per day. Starbucks pays 50 cents per muffin and charges $2.00 for it. Unsold muffins have no salvage value and cannot be carried over into the next day due to spoilage. Find the optimal stocking level. (Round to the nearest whole number)
Answer #1

Overage cost Co = Purchase price - Salvage value = 0.5 - 0 = 0.5

Underage cost Cu = Selling price - Purchase price = 2 - 0.5 = 1.5

Service level = Cu / (Cu + Co) = 1.5/(1.5+0.5) = 0.75

optimal stocking level = Minimum demand + Service level *(Maximum demand - Minimum demand)

optimal stocking level = 35 + 0.75*(80-35) = 68.75

Optimal stocking level = 68.75 Muffins